« Project Status - 19 Mar 2007 | Main | Project Status - 21 Mar 2007 »

Project Status - 20 Mar 2007

Weather Status: Storm over ARB almost a complete bust as wave basically split with energy going north and south of the experimental area. As of 20GMT Blue Canyon had no observed precipitation, despite the fact that the NAM last night was going for near an inch. Model ensembles never quit forecasting precipitation as well with amounts up to 1.5 inches even as late as the 20/06GMT runs. This mornings runs hinting at shower activity this afternoon beneath the cold core. NWS also expecting showers, so there may not be a complete shutout as the afternoon progresses. Case will be saved to study a bust situation. Beyond the end of the exercise the storm for next Mon-Tues (26-27 March) looks to be on an upward trend in the GFS and other global models. So outlook is for a possible event Monday-Tuesday. This will extend the exercise 4-5 days for ESRL staff.

System Status: Lack of rain at Slough House precluded the receiver test. All models up and running but not forecasting well.

Operations beyond 22 March: Telecons will be conducted with only ESRL personnel. Chief Scientist Fri-Tuesday: McGinley; Forecaster Fri-Sun: Szoke; Forecaster Mon-Tue: Schultz. Call participants will be all ESRL HMT staff: Kingsmill, Martner, Schneider, Neiman, White, Splitt, as available.

John McGinley, GSD

TrackBack

TrackBack URL for this entry:
/cgi-bin/mt/mt-tb.cgi/342